Two cruise ships are sailing towards each other from Caribbean islands that are 264 km apart. One ship travels 4km/h faster than the other. If they both started at the same time and meet after 6 h, how fast is each ship travelling?
(The solution is supposed to be the point of intersection between a system of two linear equations)

Answer by ankor@dixie-net.com(22740)

Two cruise ships are sailing towards each other from Caribbean islands that are 264 km apart. One ship travels 4km/h faster than the other.
If they both started at the same time and meet after 6 h, how fast is each ship travelling?
:
let x = the speed of the slower ship
then
(x+4) = the speed of the faster ship
Write a distance equation, the total distance of the two ships = 264 km
6x + 6(x+4) = 264
6s + 6s + 24 = 264
12s = 264 - 24
s = 240/12
s = 20 km/hr the slower ship, and obviously 24 km the faster
:
Check this, find the dist each travels
6(20) = 120
6(24) = 144
-----------
total d: 264
;
:
However
"(The solution is supposed to be the point of intersection between a system of two linear equations)"
let x = faster ship
let y = slower
two equations
x - y = 4
and
6x + 6y = 264
simplify, divide by 6
x + y = 44
Use elimination with these two equations
x - y = 4
x + y = 44
--------------addition eliminates y, find x
2x + 0 = 48
x = 48/2
x = 24 km/hr is the faster ship
you can find y